# From Federated Learning to $\mathbb{X}$-Learning: Breaking the Barriers of Decentrality Through Random Walks

### v1

@@ Abstract
We provide our perspective on $\mathbb{X}$-Learning ($\mathbb{X}$L), a novel distributed learning architecture that generalizes and extends the concept of decentralization. Our goal is to present a vision for $\mathbb{X}$L, introducing its unexplored design considerations and degrees of freedom. To this end, we shed light on the intuitive yet non-trivial connections between $\mathbb{X}$L, graph theory, and Markov chains. We also present a series of open research directions to stimulate further research.
1 Introduction
@@
Rapid advancement of Internet-of-Things (IoT) and widespread adoption of wireless connectivity (e.g., smart cars) have led to a surge in data-driven applications for 6G-and-beyond networks [1, 2, 3]. Collectively, modern devices generate vast amount of data, enabling the development of powerful machine learning (ML) models [4, 5]. Traditionally, these ML models are obtained using a centralized approach, requiring devices to transfer their local data to a central server. However, centralizing data can result in excessive energy consumption, high latency, and privacy concerns [6]. These challenges have fueled interest in distributed ML techniques.
1.1 Distributed ML: Federated Learning & Fog Learning
@@
Federated learning (FedL) [7, 8, 9, 10, 11, 12, 13] is the most prominent distributed ML technique that shifts model training to data-collecting devices and uses the central servers to aggregate the locally trained models of the devices. Traditional FedL begins with a central server broadcasting a global model to devices and is characterized by two sequential steps repeated until model convergence: (i) Local training, where each device updates its model using its dataset (e.g., via stochastic gradient descent (SGD)); (ii) Global model aggregations, where the devices send their models to the central server, which aggregates the received models into a new global model (e.g., via weighted averaging) and broadcasts it to all devices. Traditional FedL fits well with the device-to-edge architecture in edge networks. However, for large-scale networks (e.g., across countries/continents), FedL must be adapted to the fog network, including nodes across the cloud-to-things continuum (e.g., the cloud, core, metro, edge, and things). This has led to the development of fog learning (FogL) [14], which addresses the extreme heterogeneity in communication and computation capabilities of fog/edge network elements for effective implementation of distributed ML.
1.2 Notion of Decentralization in FedL/FogL
@@
The notion of decentralization has three major interpretations in distributed ML. (i) Centralized FedL [15, 16], where ML training is conducted at devices and the model aggregations are conducted at a central server through device-to-server communications (Fig. 1(A)). (ii) Semi-decentralized FogL/FedL [17, 18, 19, 20], which involves ML training at devices followed by decentralized model aggregations within local clusters of devices through one-hop device-to-device (D2D) communications and centralized aggregations across the clusters (Fig. 1(B)). (iii) Fully-decentralized FogL/FedL [21, 22, 23, 10], where ML training is conducted at devices and model aggregations are executed solely via one-hop D2D communications (Fig. 1(C)).
@@
1.3 From FedL/FogL to $\mathbb{X}$-Learning
@@
2 Related Work, Applications, and Challenges
2.1 Related Work
@@
(1) Gossip Learning: Gossip learning involves nodes selecting peers to exchange ML models via D2D links [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], where peer selection is often blind/random. In contrast, $\mathbb{X}$L features a purposeful/intelligent approach, where autonomous walkers/NNs traverse nodes over optimizable/deliberate trajectories and model training durations. While gossip learning is typically suited where blind/random D2D model transfers are sufficient (e.g., when nodes’ data is homogeneous), intelligent navigation across nodes combined with adaptable local ML model training make $\mathbb{X}$L applicable to heterogeneous/non-iid (non-independent and identically distributed) data across irregular network topologies.
(2) Random Walk-based Learning: Research has explored random-walk-based distributed learning, focusing on graph-specific tasks such as graph learning [37, 38, 39] (e.g., extracting structures/communities from a graph). Migrating from graph-specific tasks, $\mathbb{X}$L offers a fresh perspective to random walk-based learning by widening its scope to distributed ML tasks involving the training of single/multiple NNs over non-iid data scattered across a network of nodes. The closest work to $\mathbb{X}$L is a recent study that uses a single random walker in a graph, while balancing the exploration and exploitation of distributed data in FedL [40]. Compared to this, $\mathbb{X}$L is designed to be a generic distributed ML framework and introduces several degrees of freedom structured into two major innovations. (Innovation 1): Starting with a single walker (Sec. 3), we highlight the inadequacy of conventional graph theoretic centrality measures as well as Markov chain analysis for $\mathbb{X}$L, which triggers a set of novel concepts: (i) integration of node’ data and spatial qualities; (ii) elastic learning; (iii) perception of walkers; (iv) memory of walkers. (Innovation 2): We further envision multiple interacting walkers in $\mathbb{X}$L (Sec. 4), unveiling the following novel concepts: (i) optimal interactions/collaborations between walkers; (ii) attraction/repulsion between walkers. These innovations further trigger a variety of future directions that we will highlight.
2.2 Applications and Scenarios of Interest
@@
2.3 Challenges and Design Considerations
@@
The Vision of $\mathbb{X}$L Finding an overarching strategy addressing the above questions, combining both network- and data-dependent considerations, is the major challenge in $\mathbb{X}$L. At first glance, it may seem that answering these questions can be informed by a mixture of concepts drawn from (i) Markov chains in non-Euclidean spaces (graph topologies), (ii) graph theory, and (iii) distributed ML. However, in the subsequent discussions, we unveil that the existing frameworks/interpretations from these fields are insufficient to fully capture the dynamics inherent to $\mathbb{X}$L. This is mainly because these frameworks are developed to focus on either the graph topology (e.g., Markov chains in non-Euclidean spaces [41, 42]) or the data quality of the nodes (e.g., majority of distributed ML literature [6]). Consequently, the efficient design and implementation of $\mathbb{X}$L is an untapped area of research, which calls for a series of follow-up works. $\star$ This perspective paper does not aim to provide a formal mathematical treatment of $\mathbb{X}$L (e.g., convergence bounds or optimization formulations), which we leave as an important direction for future research. Instead, we aim to (i) unveil the fundamental concepts and degrees of freedom in $\mathbb{X}$L, (ii) articulate them intuitively, and (iii) propose simple heuristics that can achieve notable performance gains for $\mathbb{X}$L; with the ultimate goal of triggering future research.
@@
3 $\mathbb{X}$-Learning with a Single Random Walker
@@
3.1 Node Importance/Centrality
@@
In both $\mathbb{X}$L and FedL/FogL, the quality of the local data at a node is a key to determining its importance: the more representative the local data is of the global dataset (i.e., the union of the nodes’ local data), the higher its quality. Although this metric alone may suffice for FedL/FogL, it is insufficient for $\mathbb{X}$L. To demonstrate this, consider a node with poor data quality serving as a connective bridge or hub between multiple cliques of nodes, each containing high-quality data. Here, while the node itself may not offer high-quality data for training, its network centrality (i.e., spatial quality) can make it highly valuable for the walker to visit as a gateway for reaching nodes with higher-quality data. As a result, in $\mathbb{X}$L, if we aim to model the jumps/traversal between the nodes for a walker through a probability transition matrix (similar to Markov chains over graphs), this matrix should be a function of both the nodes’ data qualities and their network centralities. To achieve this goal, the concept of centrality in graph/network theory (e.g., measured via degree, betweenness, eigenvalue, and closeness metrics [45, 41, 42]) should be integrated to the concept of data quality (e.g., measured via the disparity of gradients generated on the local data of a node versus the global dataset [6]). By jointly considering spatial and data quality metrics, walkers in $\mathbb{X}$L environments follow trajectories that are adaptive to nodes’ data distribution and network topology. However, selecting a suitable realization for each metric and determining how to combine them (e.g., weighted sum of data quality and spatial quality) remains a non-trivial and open research challenge.
@@
Figure 2: Performance of random walkers under different node traversal strategies. Referring to Sec. 3.1, among methods with fixed/static node importance values (red, blue, green, magenta, and yellow curves using static sampling or fixed $\alpha\in{0,0.5,1}$), our approach with non-binary weighting of data and spatial quality metrics (yellow curve, $\alpha=0.5$) yields the best performance. Furthermore, referring to Sec. 3.3, our method with dynamic node importance (black curve using a time-varying $\alpha^{\mathsf{Inst}}$ based on current model accuracy) outperforms all the static strategies.
@@
To validate the above-described approach to node importance, we conduct a simulation and present it in Fig. 2. Here, we consider a caveman graph comprising $50$ nodes clustered into $8$ cliques. The nodes’ data distribution is non-iid with $98\%$ of nodes only containing data samples from $1$ or $2$ of the $10$ labels in CIFAR-10 dataset. We assume that the walker traverses the nodes according to various strategies, including uniform random sampling (blue curve) and Metropolis-Hastings sampling [46] (green curve) — a well-known technique in Markov chains over graphs. We consider a simple realization of our node importance approach, where the importance of node $i$ is computed as a weighted sum: $\widetilde{\mathfrak{c}}_{i}=\alpha\cdot d_{i}\cdot\ell_{i}+(1-\alpha)\cdot c_{i}$. Here, the weighting parameter $\alpha\in[0,1]$ controls the walker’s preference between data and spatial quality. In this formulation, $d_{i}$ and $\ell_{i}$ represent the fractions ($d_{i},\ell_{i}\in[0,1]$) of global dataset and labels present at node $i$ (i.e., capturing data quality), respectively, and $c_{i}$ is a conventional measure of the node’s centrality (chosen to be betweenness centrality [41, 42], capturing the spatial quality). At each node, the transition probability of the walker to the neighboring nodes is defined proportional to their importance values ($\{\widetilde{\mathfrak{c}}_{i}\}$). Our scheme (yellow curve) that obtains node importance as a balanced average ($\alpha=0.5$) between data and spatial quality, significantly outperforms other strategies, including the ones that only consider spatial quality ($\alpha=0$) and data quality ($\alpha=1$). This balanced value of $\alpha$ may not hold across all environments, but the results highlight the significance of considering network topology alongside the nodes’ data quality. We later (Sec. 3.3) show that we can further obtain dynamic transition matrices to achieve a better performance (i.e., the black curve in Fig. 2).
@@ 3.2 Elastic Learning
In Fig. 3, we show the elasticity of local training through a simulation over an RGG with $100$ nodes, $90\%$ of which have only $1$ or $2$ labels of CIFAR-10 dataset. In the three baselines, the walker is assumed to be inelastic and performing a fixed number of SGD iterations at each node ($20$, $40$, and $60$ iterations). Also, we consider an elastic walker (red curve), where its number of SGD iterations $x_{i}$ at each node $i$ is tuned according to the node’s data quality, up to a maximum of $x^{\textrm{Max}}=20$ using a sigmoid-like function $x_{i}=\frac{x^{\textrm{Max}}}{1+\exp({-\tau_{1}\times Q_{i}})}$, where $\tau_{1}=10$ is a temperature parameter, and $Q_{i}$ is a relative measure of the node $i$’s data quality defined by $Q_{i}=\ell_{i}\times{d}_{i}^{\tau_{2}\times(1-{d_{i}})}$ with $\tau_{2}=0.4$ and the rest of the variables ($d_{i}$$\ell_{i}$) are the same as Sec. 3.1. The form of this function allows for a sharp reduction in the number of SGD iterations executed at nodes with low-quality data, while its hyperparameters were chosen experimentally. The results unveil that the elastic walker performs markedly better than the inelastic ones, all while training on significantly fewer data samples. Other formulations for tuning SGD iterations according to node data quality are open for exploration.
@@
3.3 Perception-Aware Random Walk Design
We next introduce random walk perception resolution in $\mathbb{X}$L, where the walker adjusts its node traversal based on its model accuracy, shifting from static node importance-based traversals (Sec. 3.1) to accuracy-aware dynamic exploration. Specifically, early in training, with low accuracy, the walker has low perception resolution and treats most nodes as important for broad exploration. As its accuracy improves, the walker’s resolution sharpens, making it more selective, favoring nodes with high-quality data and avoiding low-quality ones to prevent performance degradation.
This further unveils a perception-aware trajectory design, where the walker’s trajectory changes over time, suggesting a departure from time-homogeneous to time-inhomogeneous Markov Chains over graphs for $\mathbb{X}$L. To show this, we redefine the node importance as a function of the walker’s model accuracy at jump/time $t$ as $\widetilde{\mathfrak{c}}_{i}(t)=\alpha^{\mathsf{Inst}}(t)\cdot d_{i}\cdot\ell_{i}+(1-\alpha^{\mathsf{Inst}}(t))\cdot c_{i}$, where compared to the fixed value of $\alpha$ used in Sec. 3.1, here $\alpha^{\mathsf{Inst}}(t)\in[\alpha^{\mathsf{Min}}{=}0.10\leavevmode\nobreak\ ,\leavevmode\nobreak\ \alpha^{\mathsf{Max}}{=}0.85]$ is a dynamic/instantaneous weighting coefficient. We model $\alpha^{\mathsf{Inst}}(t)$ as a simple function of the walker’s instantaneous accuracy $\text{A}^{\mathsf{Inst}}(t)$ as

### v2

@@ Abstract
Modern wireless networks have witnessed an unprecedented surge in data generation across distributed devices. This data can be used to empower a variety of downstream applications (e.g., autonomous driving) through machine learning (ML) techniques. Nevertheless, traditional centralized ML techniques face significant challenges in wireless networks, as they require pooling data from wireless devices at a central server, leading to high energy consumption, latency, and privacy risks. To address these issues, distributed ML techniques, such as federated learning (FedL) and fog learning (FogL), have emerged. Although these techniques enable model training directly on devices, they are often rigid in their decentralization strategies, reducing their adaptability to a wide range of wireless networks (e.g., opportunistic and volatile networks). This paper introduces $\mathbb{X}$-Learning ($\mathbb{X}$L), a novel distributed ML architecture that generalizes and extends the concept of decentralization. Inspired by web crawlers, $\mathbb{X}$L treats ML models as autonomous agents that navigate the network using a multi-hop random walk-based strategy, learning from diverse data across nodes. This flexible framework can adapt to various network conditions and topologies. Our goal is to present a vision for $\mathbb{X}$L, introducing its nuanced design considerations and unexplored degrees of freedom. To this end, we explore and unveil the foundations of $\mathbb{X}$L in both single random walker and multiple random walker settings. We also propose a set of new notions and concepts that yield novel insights on the working mechanism of $\mathbb{X}$L. We further shed light on the intuitive yet non-trivial connections between $\mathbb{X}$L, graph theory, and Markov chains. Finally, we present a series of open research directions that can trigger a variety of follow-up works.
I Introduction
@@
Rapid advancement of Internet-of-Things (IoT) and widespread adoption of wireless connectivity (e.g., smart cars) have led to a surge in data-driven applications for 6G-and-beyond networks. Collectively, modern devices generate vast amount of data, enabling the development of powerful machine learning (ML) models. Traditionally, these ML models are obtained using a centralized approach, requiring devices to transfer their local data to a central server. However, centralizing data can result in excessive energy consumption, high latency, and privacy concerns [1]. These challenges have fueled interest in distributed ML techniques.
I-A Distributed ML: Federated Learning & Fog Learning
@@
Federated learning (FedL) [2] is the most prominent distributed ML technique that shifts model training to data-collecting devices and uses the central servers to aggregate the locally trained models of the devices. Traditional FedL begins with a central server broadcasting a global model to devices and is characterized by two sequential steps repeated until model convergence: (i) Local training, where each device updates its model using its dataset (e.g., via stochastic gradient descent (SGD)); (ii) Global model aggregations, where the devices send their models to the central server, which aggregates the received models into a new global model (e.g., via weighted averaging) and broadcasts it to all devices. Traditional FedL fits well with the device-to-edge architecture in edge networks. However, for large-scale networks (e.g., across countries/continents), FedL must be adapted to the fog network, including nodes across the cloud-to-things continuum (e.g., the cloud, core, metro, edge, and things). This has led to the development of fog learning (FogL) [3], which addresses the extreme heterogeneity in communication and computation capabilities of fog/edge network elements for effective implementation of distributed ML.
I-B Notion of Decentralization in FedL/FogL
@@
The notion of decentralization has three major interpretations in distributed ML. (i) Centralized FedL, where ML training is conducted at devices and the model aggregations are conducted at a central server through device-to-server communications (Fig. 1(A)). (ii) Semi-decentralized FogL/FedL, which involves ML training at devices followed by decentralized model aggregations within local clusters of devices through one-hop device-to-device (D2D) communications and centralized aggregations across the clusters [4] (Fig. 1(B)). (iii) Fully-decentralized FogL/FedL, where ML training is conducted at devices and model aggregations are executed solely via one-hop D2D communications [5, 2] (Fig. 1(C)).
@@
I-C From FedL/FogL to $\mathbb{X}$-Learning
@@
II Related Work, Applications, and Challenges
II-A Related Work
@@
(1) Gossip Learning: Gossip learning involves nodes selecting peers to exchange ML models via D2D links [6], where peer selection is often blind/random. In contrast, $\mathbb{X}$L features a purposeful/intelligent approach, where autonomous walkers/NNs traverse nodes over optimizable/deliberate trajectories and model training durations. While gossip learning is typically suited where blind/random D2D model transfers are sufficient (e.g., when nodes’ data is homogeneous), intelligent navigation across nodes combined with adaptable local ML model training make $\mathbb{X}$L applicable to heterogeneous/non-iid (non-independent and identically distributed) data across irregular network topologies.
(2) Random Walk-based Learning: Research has explored random-walk-based distributed learning, focusing on graph-specific tasks such as graph learning [7] (e.g., extracting structures/communities from a graph). Migrating from graph-specific tasks, $\mathbb{X}$L offers a fresh perspective to random walk-based learning by widening its scope to distributed ML tasks involving the training of single/multiple NNs over non-iid data scattered across a network of nodes. The closest work to $\mathbb{X}$L is a recent study that uses a single random walker in a graph, while balancing the exploration and exploitation of distributed data in FedL [8]. Compared to this, $\mathbb{X}$L is designed to be a generic distributed ML framework and introduces several degrees of freedom structured into two major innovations. (Innovation 1): Starting with a single walker (Sec. III), we highlight the inadequacy of conventional graph theoretic centrality measures as well as Markov chain analysis for $\mathbb{X}$L, which triggers a set of novel concepts: (i) integration of node’ data and spatial qualities; (ii) elastic learning; (iii) perception of walkers; (iv) memory of walkers. (Innovation 2): We further envision multiple interacting walkers in $\mathbb{X}$L (Sec. IV), unveiling the following novel concepts: (i) optimal interactions/collaborations between walkers; (ii) attraction/repulsion between walkers. These innovations further trigger a variety of future directions that we will highlight.
II-B Applications and Scenarios of Interest
@@
II-C Challenges and Design Considerations
@@
The Vision of $\mathbb{X}$L Finding an overarching strategy addressing the above questions, combining both network- and data-dependent considerations, is the major challenge in $\mathbb{X}$L. At first glance, it may seem that answering these questions can be informed by a mixture of concepts drawn from (i) Markov chains in non-Euclidean spaces (graph topologies), (ii) graph theory, and (iii) distributed ML. However, in the subsequent discussions, we unveil that the existing frameworks/interpretations from these fields are insufficient to fully capture the dynamics inherent to $\mathbb{X}$L. This is mainly because these frameworks are developed to focus on either the graph topology (e.g., Markov chains in non-Euclidean spaces [9, 10]) or the data quality of the nodes (e.g., majority of distributed ML literature [1]). Consequently, the efficient design and implementation of $\mathbb{X}$L is an untapped area of research, which calls for a series of follow-up works. $\star$ This magazine paper does not aim to provide a formal mathematical treatment of $\mathbb{X}$L (e.g., convergence bounds or optimization formulations), which we leave as an important direction for future research. Instead, we aim to (i) unveil the fundamental concepts and degrees of freedom in $\mathbb{X}$L, (ii) articulate them intuitively, and (iii) propose simple heuristics that can achieve notable performance gains for $\mathbb{X}$L; with the ultimate goal of triggering future research.
@@
III $\mathbb{X}$-Learning with a Single Random Walker
@@
III-A Node Importance/Centrality
@@
In both $\mathbb{X}$L and FedL/FogL, the quality of the local data at a node is a key to determining its importance: the more representative the local data is of the global dataset (i.e., the union of the nodes’ local data), the higher its quality. Although this metric alone may suffice for FedL/FogL, it is insufficient for $\mathbb{X}$L. To demonstrate this, consider a node with poor data quality serving as a connective bridge or hub between multiple cliques of nodes, each containing high-quality data. Here, while the node itself may not offer high-quality data for training, its network centrality (i.e., spatial quality) can make it highly valuable for the walker to visit as a gateway for reaching nodes with higher-quality data. As a result, in $\mathbb{X}$L, if we aim to model the jumps/traversal between the nodes for a walker through a probability transition matrix (similar to Markov chains over graphs), this matrix should be a function of both the nodes’ data qualities and their network centralities. To achieve this goal, the concept of centrality in graph/network theory (e.g., measured via degree, betweenness, eigenvalue, and closeness metrics [9, 10]) should be integrated to the concept of data quality (e.g., measured via the disparity of gradients generated on the local data of a node versus the global dataset [1]). By jointly considering spatial and data quality metrics, walkers in $\mathbb{X}$L environments follow trajectories that are adaptive to nodes’ data distribution and network topology. However, selecting a suitable realization for each metric and determining how to combine them (e.g., weighted sum of data quality and spatial quality) remains a non-trivial and open research challenge.
@@
Figure 2: Performance of random walkers under different node traversal strategies. Referring to Sec. III-A, among methods with fixed/static node importance values (red, blue, green, magenta, and yellow curves using static sampling or fixed $\alpha\in{0,0.5,1}$), our approach with non-binary weighting of data and spatial quality metrics (yellow curve, $\alpha=0.5$) yields the best performance. Furthermore, referring to Sec. III-C, our method with dynamic node importance (black curve using a time-varying $\alpha^{\mathsf{Inst}}$ based on current model accuracy) outperforms all the static strategies.
@@
To validate the above-described approach to node importance, we conduct a simulation and present it in Fig. 2. Here, we consider a caveman graph comprising $50$ nodes clustered into $8$ cliques. The nodes’ data distribution is non-iid with $98\%$ of nodes only containing data samples from $1$ or $2$ of the $10$ labels in CIFAR-10 dataset. We assume that the walker traverses the nodes according to various strategies, including uniform random sampling (blue curve) and Metropolis-Hastings sampling [12] (green curve) — a well-known technique in Markov chains over graphs. We consider a simple realization of our node importance approach, where the importance of node $i$ is computed as a weighted sum: $\widetilde{\mathfrak{c}}_{i}=\alpha\cdot d_{i}\cdot\ell_{i}+(1-\alpha)\cdot c_{i}$. Here, the weighting parameter $\alpha\in[0,1]$ controls the walker’s preference between data and spatial quality. In this formulation, $d_{i}$ and $\ell_{i}$ represent the fractions ($d_{i},\ell_{i}\in[0,1]$) of global dataset and labels present at node $i$ (i.e., capturing data quality), respectively, and $c_{i}$ is a conventional measure of the node’s centrality (chosen to be betweenness centrality [9, 10], capturing the spatial quality). At each node, the transition probability of the walker to the neighboring nodes is defined proportional to their importance values ($\{\widetilde{\mathfrak{c}}_{i}\}$). Our scheme (yellow curve) that obtains node importance as a balanced average ($\alpha=0.5$) between data and spatial quality, significantly outperforms other strategies, including the ones that only consider spatial quality ($\alpha=0$) and data quality ($\alpha=1$). This balanced value of $\alpha$ may not hold across all environments, but the results highlight the significance of considering network topology alongside the nodes’ data quality. We later (Sec. III-C) show that we can further obtain dynamic transition matrices to achieve a better performance (i.e., the black curve in Fig. 2).
@@ III-B Elastic Learning
In Fig. 3, we show the elasticity of local training through a simulation over an RGG with $100$ nodes, $90\%$ of which have only $1$ or $2$ labels of CIFAR-10 dataset. In the three baselines, the walker is assumed to be inelastic and performing a fixed number of SGD iterations at each node ($20$, $40$, and $60$ iterations). Also, we consider an elastic walker (red curve), where its number of SGD iterations $x_{i}$ at each node $i$ is tuned according to the node’s data quality, up to a maximum of $x^{\textrm{Max}}=20$ using a sigmoid-like function $x_{i}=\frac{x^{\textrm{Max}}}{1+\exp({-\tau_{1}\times Q_{i}})}$, where $\tau_{1}=10$ is a temperature parameter, and $Q_{i}$ is a relative measure of the node $i$’s data quality defined by $Q_{i}=\ell_{i}\times{d}_{i}^{\tau_{2}\times(1-{d_{i}})}$ with $\tau_{2}=0.4$ and the rest of the variables ($d_{i}$$\ell_{i}$) are the same as Sec. III-A. The form of this function allows for a sharp reduction in the number of SGD iterations executed at nodes with low-quality data, while its hyperparameters were chosen experimentally. The results unveil that the elastic walker performs markedly better than the inelastic ones, all while training on significantly fewer data samples. Other formulations for tuning SGD iterations according to node data quality are open for exploration.
@@
III-C Perception-Aware Random Walk Design
We next introduce random walk perception resolution in $\mathbb{X}$L, where the walker adjusts its node traversal based on its model accuracy, shifting from static node importance-based traversals (Sec. III-A) to accuracy-aware dynamic exploration. Specifically, early in training, with low accuracy, the walker has low perception resolution and treats most nodes as important for broad exploration. As its accuracy improves, the walker’s resolution sharpens, making it more selective, favoring nodes with high-quality data and avoiding low-quality ones to prevent performance degradation.
This further unveils a perception-aware trajectory design, where the walker’s trajectory changes over time, suggesting a departure from time-homogeneous to time-inhomogeneous Markov Chains over graphs for $\mathbb{X}$L. To show this, we redefine the node importance as a function of the walker’s model accuracy at jump/time $t$ as $\widetilde{\mathfrak{c}}_{i}(t)=\alpha^{\mathsf{Inst}}(t)\cdot d_{i}\cdot\ell_{i}+(1-\alpha^{\mathsf{Inst}}(t))\cdot c_{i}$, where compared to the fixed value of $\alpha$ used in Sec. III-A, here $\alpha^{\mathsf{Inst}}(t)\in[\alpha^{\mathsf{Min}}{=}0.10~,~\alpha^{\mathsf{Max}}{=}0.85]$ is a dynamic/instantaneous weighting coefficient. We model $\alpha^{\mathsf{Inst}}(t)$ as a simple function of the walker’s instantaneous accuracy $\text{A}^{\mathsf{Inst}}(t)$ as